If sample space S={e1,e2,e3,e4}, find P(e1) if P(e2,e3)=1/5 and (e4)=1/4
1
Expert's answer
2021-07-12T13:27:16-0400
"S=({e_1,e_2,e_3,e_4}),P(e_1)=?,P(e_2,e_3)=\frac{1}{5},P({e_4})=\frac{1}{4}"

"P({e_1})+P({e_2})+P({e_3})+P({e_4})=1"

"P({e_1})+\frac{1}{2}+\frac{1}{3}+\frac{1}{4}=1"

"P({e_1})=1-(\frac{1}{2}+\frac{1}{3}+\frac{1}{4})=\frac{1}{12}"